OpenStudy (anonymous):

Look at this linear equation. 3x + 5y = 30 Which of the following equations can be used to solve for y in the linear equation above? y = - 3/5x + 30 y = 3x + 6 y = -3/5 x + 6 y = 3/5 x + 30

OpenStudy (anonymous):

can anyone show me how to do this?

OpenStudy (yanasidlinskiy):

Here how to solve the problem. You need to know the formula for this. Which is y=m(x+b) 3x+5y=30 -3x -3x 5y-3x+30 -----> Divide by 5 y= -3/5x+6

OpenStudy (the_fizicx99):

Point is to isolate y to be alone on the left hand side, Start by doing the inverse of everything, 3x + 5y = 30 you're adding 3x, so subtract 3x from both sides, 5y = -3x + 30 You're multiplying by 5, the inver is division, divide both sides by 5, 5y = -3x + 30 /5 /5 /5 y = -3/5 + 6

OpenStudy (the_fizicx99):

inverse*

OpenStudy (anonymous):

i see it now i hade to out the aquatiom into y=mx+b form?

OpenStudy (the_fizicx99):

Yes, as it's Y = mx + b